OpenStudy (anonymous):

substitution method 2x + 5y =27

OpenStudy (nothingwasthesame):

hello again

OpenStudy (k_lynn):

do you have another equation?

OpenStudy (anonymous):

so far I have.... 2( 17-3y) + 5y= 27 @(34 -6y) +5y = 27

OpenStudy (anonymous):

yea my bad...x= 17-3y

OpenStudy (k_lynn):

\[2x+5y=27\] \[x=17-3y\]

OpenStudy (k_lynn):

You need to combine the like terms